OpenStudy (anonymous):

Find (f º g)(–2) if f(x) = 4x – 3 and g(x) = 3x + 3 a. 10 b. –4 c. –15 d. 0

OpenStudy (anonymous):

That is a smart retricedog

OpenStudy (anonymous):

he ᕙ(⇀‸↼‶)ᕗ

OpenStudy (australopithecus):

sub g(x) into f(x) then set x=2

OpenStudy (australopithecus):

for example g(x) = x^2 f(x) = 3x + 1 then, (f º g)(–2) f(g(x)) = 3(x^2) + 1 f(g(-2)) = 3(-2)^2 + 1

OpenStudy (anonymous):

so d?

OpenStudy (anonymous):

@Australopithecus ?

OpenStudy (australopithecus):

not what I got

OpenStudy (australopithecus):

http://www.purplemath.com/modules/fcncomp3.htm

OpenStudy (australopithecus):

if you dont understand my solution you should check out this website it explains it pretty well

OpenStudy (australopithecus):

or example

OpenStudy (australopithecus):

you might have made an algebra mistake as well

OpenStudy (anonymous):

-4?

OpenStudy (australopithecus):

Lets go over this one step at a time what is f(g(x))?

OpenStudy (australopithecus):

hint: f(g(x)) = f(3x + 3)

OpenStudy (australopithecus):

look at my example if you are confused

OpenStudy (anonymous):

are you jake's father?

OpenStudy (australopithecus):

if you had a function f(x) = 3x we can sub stuff into the function f(2) = 3(2) f(4) = 3(4) f(-5) = 3(-5) this is the same situation with composite functions only you are subing a function into another function

OpenStudy (australopithecus):

adventure time is pretty great

OpenStudy (australopithecus):

but no that is not where my avatar is from